A two-stage rocket moves in space at a constant velocity of +4110 m/s. The two stages are then separated by a small explosive charge placed between them. Immediately after the explosion the velocity of the 1210-kg upper stage is +5540 m/s. What is the velocity (magnitude and direction) of the 2210-kg lower stage immediately after the explosion?
Answer #1

let
v = 4110 m/s
m1 = 1210, v1 = 5540 m/s
m2 = 2210 kg, v2 = ?


Apply conservation of momentum

(m1 + m2)*v = m1*v1 = m2*v2

(1210 + 2210)*4110 = 1210*5540 + 2210*v2

==> v2 = ( (1210 + 2210)*4110 - 1210*5540 )/2210

= 3327 m/s <<<<<<<<------------------------------------------Answer

direction : in the same direction